OpenStudy (anonymous):

parallel and perpendicular equations for the point (-5,-10) 2x + 5y -12 = 0

OpenStudy (amistre64):

parallel have same equation with a difference constat; perps switch up the xy coeffs and negate one; and change the constant too

OpenStudy (amistre64):

// 2x + 5y + c = 0 ; calibrate with the given point L 5x - 2y + c = 0 ; calibrate with given point

OpenStudy (anonymous):

so if you calculate it out as in slope intercept form, and got y=-2/5 = 12/5 perpendicular would just be 5/2+ whatever else?

OpenStudy (anonymous):

that was supposed to be + 12/5

OpenStudy (amistre64):

looks to be a good set up yes

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

slope is -2/5 for parallel \\ slope is 5/2 for perp L

OpenStudy (amistre64):

then ignore the 12 and use the point fo find a new constant

OpenStudy (amistre64):

y = -2/5 x + c -10 = -5(-2/5) + c -10 = 2 + c -12 = c y = -2/5 - 12 ..... that worked out kinda cool lol

OpenStudy (amistre64):

forgot the "x" in there tho ... bad amistre, bad!! :)

OpenStudy (anonymous):

lol thank you:)

OpenStudy (amistre64):

yw :) the perp is the same concept; ad it looks like your competent enough for that one

OpenStudy (anonymous):

I know math, I just doubt myself. Whenever I get my work back it's always STUPID little mistakes